OpenStudy (anonymous):

Simplify the following expression. 6^3 x 6^-5

jagr2713 (jagr2713):

Well first we simplify 6^3 to 216 to get \[216x \times \frac{ 1 }{ 6^{5} }\]

jagr2713 (jagr2713):

Opps no its \[216x \times 6^{-5}\] Disregard the first equation

jagr2713 (jagr2713):

Next we use the negative power rule :\[x ^{-a}=\frac{ 1 }{ x ^{a} }\] what o we get when we use that rule

OpenStudy (anonymous):

uhm.. i dunno...

jagr2713 (jagr2713):

Well i said it above by accident.

OpenStudy (anonymous):

xD urm, negetave numbers? Lool i feel so stupid im horrible with math. Ermm, Yeah, Negetave powers?

jagr2713 (jagr2713):

A negative exponent just means that the base is on the wrong side of the fraction line, so you need to flip the base to the other side. For instance, "x–2" (ecks to the minus two) just means "x2, but underneath, as in 1/(x2)". So we would get \[216x \times \frac{ 1 }{ 6^{5} }\]

jagr2713 (jagr2713):

So from that what do we simplify?

OpenStudy (anonymous):

1/6^5

jagr2713 (jagr2713):

Nope whats \[6^{5}\]

jagr2713 (jagr2713):

You there @freakyaoifrangirl